1.Jenny wants to utilize a portion of her garden for growing flowers. Being a designer herself, she visualizes a triangular area with different lengths for each side. For this Jenny decides to keep a side of the triangular area 3 feet shorter and the other 2 feet longer than the third side. She also wants to restrict the perimeter of the triangular area to 32 feet, so that it does not cover a huge area of the garden. Determine the maximum length of each side of the triangular area that Jenny has visualized.
Answer by stanbon(75887)

Jenny wants to utilize a portion of her garden for growing flowers. Being a designer herself, she visualizes a triangular area with different lengths for each side. For this Jenny decides to keep a side of the triangular area 3 feet shorter and the other 2 feet longer than the third side. She also wants to restrict the perimeter of the triangular area to 32 feet, so that it does not cover a huge area of the garden. Determine the maximum length of each side of the triangular area that Jenny has visualized.
-----------------
Letl length of the 3rd side be "x" ft.
Then length of the 2nd side is "x-3"
And length of the 1st side is "x+2"
--------------------------------------------
Equation:
Perimeter = 32
x + x+2 + x-3 = 32
------
3x 33
x = 11 ft.(3rd side)
x-3 = 8 ft
x+2 = 13 ft
====================
Cheers,
Stan H.